OpenStudy (anonymous):

Calculate the number of molecules in 6.00 moles h2s?

OpenStudy (aaronq):

multiply the number of moles by avogadros number. \(n=\dfrac{N}{N_A}\rightarrow N=n*N_A\) n=moles; N=number of particles (in this case, molecules); \(N_A\)=avogadro's number